1.what is the wavelength associated of an electron moving from n=4 state to n=2 state in a hydrogen atom ( Explain in detail, please)
Answer #1

Here photon will be emitted

1/lambda = R* (1/nf^2 - 1/ni^2)

R is Rydberg constant. R = 1.097*10^7 m-1

1/lambda = R* (1/nf^2 - 1/ni^2)

1/lambda = 1.097*10^7* (1/2^2 - 1/4^2)

lambda = 4.862*10^-7 m

lambda = 486 nm

Answer: 486 nm